# Large Language Model Guided Self-Debugging Code Generation

### v1

@@ Abstract.
Automated code generation is gaining significant importance in intelligent computer programming and system deployment. However, current approaches often face challenges in computational efficiency and lack robust mechanisms for code parsing and error correction. In this work, we propose a novel framework, PyCapsule, with a simple yet effective two-agent pipeline and efficient self-debugging modules for Python code generation. PyCapsule features sophisticated prompt inference, iterative error handling, and case testing, ensuring high generation stability, safety, and correctness. Empirically, PyCapsule achieves up to 5.7% improvement of success rate on HumanEval, 10.3% on HumanEval-ET, and 24.4% on BigCodeBench compared to the state-of-art methods. We also observe a decrease in normalized success rate given more self-debugging attempts, potentially affected by limited and noisy error feedback in retention. PyCapsule demonstrates broader impacts on advancing lightweight and efficient code generation for artificial intelligence systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Software and its engineering Automatic programmingccs: Computing methodologies Natural language processing
@@ 1. Introduction
The evolution of generative artificial intelligence has profoundly influenced multiple domains, notably in automated code generation. Particularly, Large Language Models (LLMs) have demonstrated remarkable proficiency in code generation given natural language prompts, streamlining software development processes and enhancing productivity  (Rasnayaka et al., 2024; Liu et al., 2024; Chowdhery et al., 2024). This transformative potential has sparked significant research in this field. However, several critical challenges emerge as these systems scale to handle increasingly complex programming tasks. Ensuring reliability, accuracy, and autonomy in code generation remains critical but challenging, particularly in error detection and debugging, that align with human developers’ intentions  (Fan et al., 2023; Dou et al., 2024; Lian et al., 2024). Recent research has introduced frameworks consisting of multiple LLM-based agents to emulate human problem-solving capability in code generation. For instance, MapCoder  (Islam et al., 2024), AgentCoder  (Huang et al., 2024b), and LDB  (Zhong et al., 2024) employ agents dedicated to retrieval, planning, coding, and debugging to enhance code quality. While these approaches have shown remarkable improvements, they often require substantial computational resources and complex agent coordination. This poses challenges for real-world deployment and scalability.
@@
To address these concerns, we propose a novel framework, PyCapsule, that enhances the efficiency and accuracy of LLM-based Python code generation through a two-agent architecture. This simple yet effective two-agent interaction ensures functionality independence without overcrowding with AI agents, thereby minimizing its computational consumption. PyCapsule employs a programmer agent responsible for code generation and debugging, interacting with an executor agent that handles code validation, case testing, and error analysis. The framework’s effectiveness is further enhanced by three specialized modules: an error handler that refines the debugging feedback, an example call detector that prevents runtime issues, and a function signature converter that provides a uniform template for the function head description. This structure enables PyCapsule to achieve high performance on widely used datasets, including HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), Mostly Basic Python Programming (MBPP) (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench  (Zhuo et al., 2024) with a high computational efficiency. The main contributions of our work are summarized as
@@
  • Our PyCapsule is a resource-efficient framework that achieves high performance with significantly reduced computational overhead compared to existing multi-agent approaches. Our two-agent architecture demonstrates that complex coordination between multiple agents can be replaced by well-structured language prompts and modules without sacrificing accuracy.
@@
  • Our efficient debugging mechanism aids the programmer agent through three specialized modules: an error handler, an example call detector, and a function signature converter. These achieve up to 5% accuracy improvement on HumanEval, 10% on HumanEval-ET using GPT models, and 25% on BigCodeBench using the Qwen 7B model compared to its 32B variant.
@@
  • Through comprehensive analysis across multiple datasets and models, we observe a consistent decrease in the normalized success rate as we increase the number of self-debugging attempts. We hypothesize that this is caused by the increased problem complexity due to limited and noisy error messages for the self-debugging process.
@@ 2. Related Work
Automated high-quality code generation has evolved in several key directions related to prompt engineering (Knoth et al., 2024; Wei et al., 2022), multi-agent systems (Wooldridge and Jennings, 1995; Burgin and Dodig Crnkovic, 2009), and iterative debugging with runtime feedback mechanisms (Park et al., 2023; Chen et al., 2023; Islam et al., 2024; Huang et al., 2024b; Li et al., 2025). While LLM has demonstrated increasing capability in code generation, the code quality and correctness can be further enhanced by using intelligent debugging strategies. Therefore, previous works in self-debugging (Huang et al., 2023; Chen et al., 2023) show their capability in Automated Program Repair (APR) by identifying and fixing significant code flaws and bugs. Researchers have shown that ”large language models of code” (Huang et al., 2023) often struggle to directly fix bugs in the generated code due to the lack of task-specific fine-tuning. These fine-tuned models outperform the state-of-the-art APR tools (Xia and Zhang, 2022b; Fu et al., 2022; Joshi et al., 2023) but still suffer from the loss of pre-trained knowledge (Xia and Zhang, 2022a; Adnan et al., 2024), lack of debugging ingredients, inferior long-sequence handling ability, high resource constraints, long-tail problems, and multi-hunk error fixing (Huang et al., 2023).
@@
Particulary, MapCoder (Islam et al., 2024) presents a multi-agent framework designed to emulate human-level programming processes. It employs four agents for retrieval, planning, coding, and debugging to perform iterative code generation and debugging. AgentCoder (Huang et al., 2024b) introduces a three-agent framework to address high token usage and coordination overhead observed in other multi-agent systems like MapCoder (Islam et al., 2024), MetaGPT (Hong et al., 2024), and ChatDev (Qian et al., 2024). The framework comprises a programmer agent, a test designer agent, and a test executor agent. The ”Debug Like a Human” framework (Zhong et al., 2024) introduces Large Language Model Debugger (LDB), which enhances the code quality by leveraging runtime execution information. LDB is performed iteratively through segmenting code, tracking intermediate variable states, and performing step-by-step verification with breakpoints. Moreover, some approaches have employed prompt engineering to improve the quality of code generation and debugging. For instance, CodeCoT (Huang et al., 2024a) leverages iterative refinement through Chain of Thought (CoT) (Wei et al., 2022) and logical reasoning. However, its capability is limited to refining syntax errors.
@@
While these approaches have advanced automated code generations, they share several notable limitations, particularly the resource intensity for complex problems. Additionally, they often struggle to fully leverage error messages, which impairs their adaptability to nuanced debugging scenarios. Although this issue can be alleviated by generating complementary test cases, it can further increase resource consumption due to the reliance on preemptive test cases such that generation may become less reliable  (Li et al., 2025; Li and Yuan, 2024) and potentially require more complex agents.
@@
To address these challenges, we introduce PyCapsule as a streamlined modular architecture with robust and optimized modules. It combines the knowledge from the operation research field and computer science to optimize the processing pipeline with merely two AI agents. This greatly reduces the computational overhead and facilitates more efficient and reliable code generation.
@@ 3.1. Programmer and Executor Agents
Figure 1. PyCapsule framework as Python code generation service of OpenSI-CoSMIC (Adnan et al., 2024). PyCapsule consists of iterative code generation with debugging, error handling, and code execution within a Docker container.
@@
PyCapsule employs two distinct types of agents as defined in the agent-based computing literature (Nwana, 1996; Jennings et al., 1998): smart agents and collaborative agents. The system consists of two agents: a programmer agent and an executor agent. The programmer agent generates code and handles debugging while the executor agent acts as an autonomous validator that continuously monitors and responds to code execution outcomes. When execution succeeds, it forwards the validated code to the user; otherwise, it generates detailed diagnostic reports and initiates a real-time feedback loop with the programmer agent for self-debugging.
@@
The programmer agent leverages Ollama and GPT infrastructure embedded in OpenSI-CoSMIC (Adnan et al., 2024) to facilitate self-debugging and code generation through a structured pipeline. The programmer agent employs tailored system prompts with two programming modes, generation mode and fix mode. These prompts, detailed in Appendix  E, define the programmer agent’s persona (Shanahan et al., 2023), provide task-specific background, and structurally simplify code extraction. In the generation mode, the programmer agent uses CoT (Wei et al., 2022) to analyze the problem description and devise a structured solution to generate executable code. In the fix mode, the executor agent uses error information from the previous debugging attempt. This mode is supported by the original problem description, prior solution, and error messages processed by the error-handling module. The conversation history includes the most recent problem-solution pair. For MBPP with Qwen2.5-Coder-7B-Instruct, tests with 2 and 3 conversation pairs decrease the accuracy from 80.7% to 77.6% and 76.7%, respectively.
We implemented the executor agent using a lightweight Docker container. The container’s entry point is a shell script for installing required Python libraries and testing the generated code, which are generated by the programmer agent. Once the problem is solved, the container returns a success or failure status. The success status indicates that the generated code is correct by passing all the test cases while the failure status will trigger the error-handling module to refine and send the error message to the programmer agent. The programmer agent will then enter the fix mode to fix code flaws iteratively with self-debugging attempts. Particularly, with our analysis in Section 4.2 showing the diminishing influence of successive debugging attempts, the programmer agent is capable of refining the incorrect code with up to 5 self-debugging attempts. This seamless interaction between the two agents ensures robust and efficient feedback-driven debugging.
@@ 3.2. Supportive Modules
The signature converter mitigates the instability of code generation by parsing problem descriptions provided in datasets through inferring structured function signatures. This module generates a function name, a robust signature, and an example call using the first test case for a given problem without revealing the case-testing result, thereby avoiding extra LLM calls for signature inference.
@@
The error-handling module refines error messages provided by the executor agent, giving concise and highly relevant natural language feedback. It also identifies the error type, filters irrelevant tracebacks, and resolves critical issues such as verbose recursion errors that can disrupt the programmer agent due to the LLM’s limited context length. This systematic approach requires fewer tokens in self-debugging iterations compared to the raw but redundant error messages.
These modules collectively enhance the code generation stability, safety, and self-debugging efficiency while reducing the computational overhead. More details are provided in Appendix C.
@@ 3.3. Code Generation Workflow
PyCapsule consists of three phases: code generation, execution, and self-debugging, shown in Figure 1. This pipeline integrates the programmer agent and executor agent with the self-bugging mechanism to ensure reliable code generation and case testing with high efficiency and accuracy.
The workflow starts with the programmer agent that receives a problem description, typically supplemented by a function signature either from the dataset or the signature converter module. Subsequently, the agent generates a response with CoT reasoning, environment setup instructions, and code synthesis. Then, the agent’s response is parsed to extract the synthesized code and environment setup instructions, with the latter automatically generating a text file, requirements.txt, containing all required Python libraries.
The extracted code is further processed through the example call detector to remove any unintended example calls. At the generation phase, the processed code with associated test cases from the dataset is saved in a Python entry file main.py for code execution. To ensure code execution safety, we add a timer with a maximum execution time to prevent infinite loops and raise an out-of-time error for those requiring longer running time.
At the execution phase, the executor agent runs main.py in a Docker container with Python libraries from requirements.txt installed. The corresponding code execution status, either successful or failed, determines whether the current generated code requires to be fixed by using the programmer agent’s fix mode. If so, error messages extracted from the executor agent are refined by the error-handling module and will used by the programmer agent to fix code flaws. This mode allows up to 5 self-debugging attempts, each incorporating code execution feedback from the previous bug-fixing process. The pseudo-code for the pipeline is provided in Appendix D.
@@ 4. Experiments
All GPU-related experiments were conducted on a single NVIDIA GeForce RTX 3090 (24 GB). We evaluated performance using success rate, defined by the ratio of correctly resolved problems to total problems. Our code will be released upon publication.
@@ 4.1. Dataset Overview
We used five benchmark datasets to evaluate PyCapsule: HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), MBPP (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench (Zhuo et al., 2024), where the ”ET” variants provide more test cases and enriched problem descriptions. These datasets encompass diverse programming challenges.
Particularly, HumanEval (Chen et al., 2021) consists of 164 human-curated programming problems for string manipulation, arithmetic, and basic data structures. Each problem includes a natural language description, a function signature, and test cases. MBPP (Austin et al., 2021) includes 974 programming problems spanning algorithms, data structures, and general-purpose tasks. It provides natural language descriptions and test cases but no function signatures to guarantee code generation stability. BigCodeBench (Zhuo et al., 2024) includes 1,140 problems with two splits: the complete split for code completion with specific docstrings and the instruct split with concise instructions for advanced reasoning.

### v2

@@ Abstract.
Automated code generation is gaining significant importance in intelligent computer programming and system deployment. However, current approaches often face challenges in computational efficiency and lack robust mechanisms for code parsing and error correction. In this work, we propose a novel framework, PyCapsule, with a simple yet effective two-agent pipeline and efficient self-debugging modules for Python code generation. PyCapsule features sophisticated prompt inference, iterative error handling, and case testing, ensuring high generation stability, safety, and correctness. Empirically, PyCapsule achieves up to 5.7% improvement of success rate on HumanEval, 10.3% on HumanEval-ET, and 24.4% on BigCodeBench compared to the state-of-art methods. We also observe a decrease in normalized effectiveness given more self-debugging attempts, potentially affected by limited and noisy error feedback in retention. PyCapsule demonstrates broader impacts on advancing lightweight and efficient code generation for artificial intelligence systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Automatic programmingccs: Computing methodologies Natural language processing
@@ 1. Introduction
The evolution of generative artificial intelligence has profoundly influenced multiple domains, notably in automated code generation. Particularly, Large Language Models (LLMs) have demonstrated remarkable proficiency in code generation given natural language prompts, streamlining software development processes and enhancing productivity  (Rasnayaka et al., 2024; Liu et al., 2024; Chowdhery et al., 2024). This transformative potential has sparked significant research in this field. However, several critical challenges emerge as these systems scale to handle increasingly complex programming tasks. Ensuring reliability, accuracy, and autonomy in code generation remains critical but challenging, particularly in error detection and debugging, that aligns with human developers’ intentions  (Fan et al., 2023; Dou et al., 2024; Lian et al., 2024). Recent research has introduced frameworks consisting of multiple LLM-based agents to emulate human problem-solving approaches in code generation. For instance, MapCoder  (Islam et al., 2024), AgentCoder  (Huang et al., 2024b), and LDB  (Zhong et al., 2024) employ agents dedicated to retrieval, planning, coding, and debugging to enhance code quality. While these approaches have shown remarkable improvements, they often require substantial computational resources and complex agent coordination. This poses challenges for real-world deployment and scalability.
@@
To address these concerns, we propose a novel framework, PyCapsule, that enhances the efficiency and accuracy of LLM-based Python code generation through a two-agent architecture. This simple yet effective two-agent interaction ensures functionality independence without overcrowding with AI agents, thereby minimizing its computational consumption. PyCapsule employs a programmer agent responsible for code generation and debugging, interacting with an executor agent that handles code validation, case testing, and error analysis. The framework’s effectiveness is enhanced by three specialized modules: an error handler that refines the debugging feedback, an example call detector that prevents runtime issues, and a function signature converter that provides a uniform template for the function description. This structure enables PyCapsule to achieve high performance on widely used datasets, including HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), Mostly Basic Python Programming (MBPP) (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench (full) (Zhuo et al., 2024) with a high computational efficiency. The main contributions of our work are summarized as -
@@
  • Our two-agent architecture implemented in PyCapsule demonstrates that complex coordination between multiple agents can be replaced by well-structured language prompts and modules without sacrificing accuracy.
@@
  • We introduce an efficient debugging mechanism by combining the programmer agent and executor agent with three specialized modules: an error handler, an example call detector, and a function signature converter. These achieve up to 5% accuracy improvement on HumanEval, 10% on HumanEval-ET using GPT-3.5-Turbo-1106 and GPT-4-Preview-1106, and 25% on BigCodeBench using the Qwen-2.5-coder-instruct 7B model.
@@
  • Through comprehensive analysis across multiple datasets and models, we demonstrate that there is a consistent decrease in the normalized debugging success rate as the number of self-debugging attempts increases. This suggests that the growing complexity of problems, exacerbated by limited and noisy error messages, may be hindering the self-debugging process.
@@ 2. Related Work
Automated high-quality code generation has evolved in several key directions related to prompt engineering (Knoth et al., 2024; Wei et al., 2022), multi-agent systems (Wooldridge and Jennings, 1995; Burgin and Dodig Crnkovic, 2009), and iterative debugging with runtime feedback mechanisms (Park et al., 2023; Chen et al., 2023; Islam et al., 2024; Huang et al., 2024b; Li et al., 2025a). While LLM has demonstrated increasing capability in code generation, the code quality and correctness can be further enhanced by using intelligent debugging strategies. Therefore, previous works in self-debugging (Huang et al., 2023; Chen et al., 2023) show their capability in Automated Program Repair (APR) by identifying and fixing significant code flaws and bugs. Researchers have shown that ”large language models of code” (Huang et al., 2023) often struggle to directly fix bugs in the generated code due to the lack of task-specific fine-tuning. These fine-tuned models outperform the state-of-the-art APR tools (Xia and Zhang, 2022b; Fu et al., 2022; Joshi et al., 2023) but still suffer from the loss of pre-trained knowledge (Xia and Zhang, 2022a), lack of debugging ingredients, inferior long-sequence handling ability, high resource constraints, long-tail problems, and multi-hunk error fixing (Huang et al., 2023).
@@
Particulary, MapCoder (Islam et al., 2024) presents a multi-agent framework designed to emulate human-like programming practices. It employs four agents for retrieval, planning, coding, and debugging to perform iterative code generation and debugging. AgentCoder (Huang et al., 2024b) introduces a three-agent framework to address high token usage and coordination overhead observed in other multi-agent systems like MapCoder (Islam et al., 2024), MetaGPT (Hong et al., 2024), and ChatDev (Qian et al., 2024). The framework comprises a programmer agent, a test designer agent, and a test executor agent. The ”Debug Like a Human” framework (Zhong et al., 2024) introduces Large Language Model Debugger (LDB), which enhances the code quality by leveraging runtime execution information. LDB is performed iteratively through segmenting code, tracking intermediate variable states, and performing step-by-step verification like breakpoints. Moreover, some approaches have employed prompt engineering to improve the quality of code generation and debugging. For instance, CodeCoT (Huang et al., 2024a) leverages iterative refinement through Chain of Thought (CoT) (Wei et al., 2022) and logical reasoning. However, its capability is limited to refining syntax errors.
@@
While these approaches have advanced automated code generations, they share several notable limitations, particularly the resource intensity for complex problems. Additionally, they often struggle to fully leverage error messages, which impairs their adaptability to nuanced debugging scenarios. Although this issue can be alleviated by generating complementary test cases, this approach increases computational demands and may reduce reliability (Li et al., 2025a; Li and Yuan, 2024). Furthermore, this test-based compensation strategy potentially necessitates more sophisticated agent architectures.
@@
To address these challenges, we introduce PyCapsule as a streamlined modular architecture with robust and optimized modules. It combines the knowledge from the operation research field and computer science to optimize the processing pipeline with merely two AI agents. This greatly reduces the framework complexity and facilitates more efficient and reliable code generation.
@@ 3.1. Programmer and Executor Agents
Figure 1. PyCapsule framework for Python code generation. The framework comprises iterative code generation, debugging, error handling, and code execution within a Docker container.
@@
PyCapsule employs two distinct types of agents, as defined in the agent-based computing literature (Nwana, 1996; Jennings et al., 1998): smart agents and collaborative agents. The system includes two such agents — a programmer agent and an executor agent — which embody these types, respectively. Following the agent typology in (Nwana, 1996), these agents are classified based on their emphasis on three key attributes: autonomy, cooperation, and learning. The programmer agent functions as a smart agent, emphasizing all three attributes through its LLM-based reasoning, autonomous code generation decisions, cooperative debugging interactions, and performance improvements over debugging iterations. The executor agent operates as a collaborative agent, which, while implemented with a Docker container and supporting modules rather than an LLM, meets the essential criteria of autonomy (operating without direct human intervention) and cooperation (communicating execution results and coordinating with the programmer agent). The executor agent acts as an autonomous validator that continuously monitors and responds to code execution outcomes, makes independent decisions about success or failure, and cooperates with the programmer agent by providing structured feedback to facilitate debugging. When execution succeeds, it forwards the validated code to the user; otherwise, it generates detailed diagnostic reports and initiates a real-time feedback loop with the programmer agent for self-debugging. This interaction exemplifies the social ability identified in (Jennings et al., 1998) as necessary for agent cooperation.
@@
The programmer agent employs tailored system prompts with two programming modes, generation mode and fix mode. These prompts, detailed in Appendix  F, define the programmer agent’s persona (Shanahan et al., 2023), provide task-specific background, and simplify code extraction (Li et al., 2025b). In the generation mode, the programmer agent uses Chain-of-Thought(CoT) (Wei et al., 2022) to analyze the problem description and devise a structured solution to generate executable code. In the fix mode, the programmer agent attempts to debug code using error information from the previous execution attempt. This fix mode is supported by three key information elements: the original problem description, the prior solution attempt, and error messages processed by the error-handling module. For conversation history management, we retain only the most recent problem-solution pair, consisting of the original problem description, prior solution attempt, and processed error messages—as this single pair defines the context for the current debugging attempt. The choice to retain only the most recent aligns with Markov Decision Process (MDP), where the agent’s decision depends only on the current state (Sutton and Barto, 2018). Furthermore, this design addresses a known challenge of LLM performance degradation with increasing context length—a phenomenon documented in recent literature on attention dilution and information retrieval difficulties in longer contexts (Li et al., 2025b). Our empirical testing confirms this effect: when evaluating MBPP with Qwen2.5-Coder-7B-Instruct, increasing from one to two or three conversation pairs progressively degraded performance (80.7% → 77.6% → 76.7%), supporting our decision to maintain minimal conversation history.
We implemented the executor agent with access to a lightweight Docker container. The container uses a predefined shell script as its entry point, which handles environment setup and code execution. While this infrastructure script is part of the executor agent and static, the programmer agent generates the solution code and any required dependency specifications for each problem. This separation allows for a consistent execution environment while enabling problem-specific implementations. Once the problem is solved, the container returns either a success or a failure status. The success status indicates that the generated code has passed all test cases. At the same time, failure triggers the error-handling module to relay diagnostics to the programmer agent, which then enters fix mode for iterative self-debugging. Based on preliminary experiments that showed diminishing returns for additional debugging iterations, the programmer agent is limited to a maximum of five self-debugging attempts. This limit was empirically determined through exponential decay analysis of debugging effectiveness (see 4.2). Each attempt incorporates code execution feedback from the previous debugging attempt.
@@ 3.2. Supportive Modules
The signature converter enhances code generation stability by extracting structured function signatures from problem descriptions. This module generates a function name, a robust signature, and an example call using the first test case for a given problem without revealing the test results, thereby avoiding extra LLM calls for signature inference.
@@
The error-handling module refines error messages provided by the executor agent, giving concise and highly relevant natural language feedback. It also identifies the error type, filters irrelevant tracebacks, and resolves critical issues such as verbose recursion errors that can disrupt the programmer agent due to the LLM’s limited context length. This systematic approach limits token usage during self-debugging iterations compared to the raw but redundant error messages.
These modules collectively enhance the code generation stability, safety, and self-debugging efficiency while reducing the computational overhead. More details are provided in Appendix E.
@@ 3.3. Code Generation Workflow
PyCapsule consists of three phases: code generation, execution, and self-debugging, shown in Figure 1. This pipeline integrates the programmer and executor agents with the self-debugging mechanism to ensure reliable code generation, high efficiency, and accuracy in testing.
The workflow starts with the programmer agent receiving a problem description, typically supplemented by a function signature from the dataset or the signature converter module. Subsequently, the agent generates a response with CoT reasoning, environment setup instructions, and code synthesis. This response is parsed to extract the synthesized code and environment setup instructions, with the latter automatically generating a text file, requirements.txt, containing all required Python libraries.
The extracted code is further processed through the example call detector to remove unintended example calls. The final processed code and associated test cases from the dataset are saved in a Python file main.py for code execution. For execution safety, we implement a maximum runtime timer to prevent infinite loops and handle timeout errors.
At the execution phase, the executor agent sets the Python environment using the requirements from requirements.txt and runs main.py in a Docker container. The corresponding code execution status, either successful or failed, determines whether the current generated code needs to be fixed by using the programmer agent’s fix mode. If so, error messages extracted from the executor agent are refined by the error-handling module and will be used by the programmer agent to fix code flaws. This mode allows up to five self-debugging attempts, each incorporating code execution feedback from the previous bug-fixing process. The pseudo-code for the pipeline is provided in Appendix D.
@@ 4. Experiments
All GPU-related experiments were conducted on a single NVIDIA GeForce RTX 3090 (24 GB). We evaluated performance using success rate, defined by the ratio of correctly resolved and total problems. Our code will be released upon publication.
@@ 4.1. Dataset Overview
We used five benchmarks to evaluate PyCapsule: HumanEval (Chen et al., 2021), HumanEval-ET (Dong et al., 2023), MBPP (Austin et al., 2021), MBPP-ET (Dong et al., 2023), and BigCodeBench(full) (Zhuo et al., 2024), where the ”ET” variants provide more test cases and enriched problem descriptions. These datasets encompass diverse programming challenges.
Particularly, HumanEval (Chen et al., 2021) consists of 164 human-curated programming problems for string manipulation, arithmetic, and basic data structures. Each problem includes a natural language description, a function signature, and test cases. MBPP (Austin et al., 2021) includes 974 programming problems spanning algorithms, data structures, and general-purpose tasks. It provides natural language descriptions and test cases, but no function signatures to guarantee code generation stability. BigCodeBench(full) (Zhuo et al., 2024) includes 1,140 problems with two prompt splits: the complete split for code completion with specific docstrings and the instruct split with concise instructions for advanced reasoning. We evaluated our system using both splits.